OpenStudy (anonymous):

Find the equations of the vertical and horizontal asymtotes. If an asymptote does not exist, state "none". f(x) = (x^3)/(x^2+1) VA=? HA=?

OpenStudy (goformit100):

diffretiate it and put the values in it.

OpenStudy (anonymous):

Huh??? You lost me.

OpenStudy (anonymous):

HA = (x^3)/(x^2) = ? VA = x^2+1 = ?

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

OK so I am getting no HA or VA is that correct?

OpenStudy (amistre64):

it has 2 VA (what makes the bottom of it go zero that doesnt factor with the top) and no HA (compare leading terms x^3/x^2 = x .. not a constant, or a 1/x deal.)